Kim Wilson borrowed $7,800 from Guaranteed Bank for 24 months. Assuming the bank charges Kim 7.625% annually for this loan, calculate the following. (Round your responses to the nearest cent if necessary.)

A) Interest?

B) Maturity?
Answer #1
Statement showing Computations
Particulars Amount
A= P(1+r/100)^n
A= 7800 *(1+.07625)^2
A= 7800 *(1.07625)^2
A= $9,034.85
Interest = 9034.85 - 7800                  1,234.85
Maturity =                    9,034.85